# Look as You Leap: Planning Simultaneous Motion and Perception for High-d o f Robots

### v1

@@ IV-A Neural Surrogate Perception Score Models
To address these challenges, we propose a neural surrogate model inspired by [80, 81] that approximates perception scores without requiring explicit rendering. Alg. 2 outlines the full pipeline. Given a robot configuration and the target object label, the surrogate model predicts the expected perception performance (e.g.,, detection confidence for a cup). We first apply FK to map the configuration from $\mathcal{Q}$ to the robot’s camera pose. The neural model is then trained to map these camera transforms to perception scores obtained from labeled, photorealistic simulations. Specifically, we generate a dataset by rendering views of objects or humans from varied camera poses in simulation and evaluating each image with a perception model. The surrogate is implemented as a multilayer perceptron (MLP) [82] that takes the relative pose between camera and object, along with an object encoding, and outputs the corresponding perception score. Trained on this dataset, the surrogate enables fast and scalable evaluation during planning. Importantly, because the input is based on camera pose rather than robot-specific joint values, the model generalizes across robots with similar perception setups.
To address object inter-occlusion in cluttered indoor environments, we build upon techniques similar to those in [83, 84] by performing ray-casting towards surrounding objects. This step determines whether the object of interest is occluded by checking for intersections between each cast ray and the Axis-Aligned Bounding Boxes (AABB) of nearby objects. For a new robot configuration, we obtain an occlusion score in addition to the perception score estimated by the neural surrogate model. This occlusion score is computed as the ratio of rays that ultimately do not intersect the bounding box of the target object—–an indicator that something else is blocking the view. If the occlusion score exceeds a chosen threshold, we set the perception score of the neural surrogate model to zero for that configuration. Fig. 4 shows a detailed overview of this occlusion checking step.
@@ IV-B1 Sampling with Visibility Constraints
For high-d o f robots, computing motion that considers the aforementioned object-centered models can be achieved using sampling-based motion planning (SBMP) [86] with manifold constraints [87] where planning happens on a subspace of the free configuration space defined by a manifold. Specifically, in the construction of our proposed planner, As shown in Alg. 1, ps-prm first constructs a PRM with states and transitions that satisfy the manifold constraints, using either line-ofsight or camera frustum. The neighbor nodes in the PRM are constrained to lie on the same manifold and are selected based on their geodesic proximity within a certain distance. To sample from the manifold, ps-prm first selects a robot configuration from $\mathcal{Q}_{free}$ and then applies the projection operator to map it onto the surface of $M$. This involves iteratively retracting the point towards a minimum of the constraint function and solving a linear system of equations at each step. Configurations that cannot be projected into valid in-manifold configurations are discarded, and the algorithm resamples until a valid configuration is found or a timeout is reached. This process ensures that the object is placed either within the line of sight or inside the camera frustum.
@@ IV-B2 PRM construction and scoring
After the PRM is constructed and scored, ps-prm connects the start and goal to the PRM, and uses A* [88] with an admissible heuristic to find a path between the start and goal configurations. In this paper, we demonstrate experiments on object detection and tracking, as well as human face detection, though the approach can be generalized to other perception tasks.
@@ IV-C GPU-Parallelized Replanning Pipeline
The proposed perception scoring methods are highly parallelizable. Since the perception scores are computed using a neural surrogate model specifically optimized for GPU inference, we are motivated to implement the entire planning pipeline on the GPU. This GPU-accelerated pipeline performs collision checking, constraint projection, FK, and model inference entirely in parallel, significantly improving runtime efficiency. PS-PRM can therefore operate effectively in dynamic environments where the object of interest may move during execution. In such cases, the planner updates node perception scores in real time and replans the path as needed to adapt to the new configuration. Moreover, communication between the GPU-equipped machine and the robot is lightweight, as only the planned trajectories are transmitted. Fig. 5 shows a detailed overview of the GPU-accelerated process.
When the object of interest moves within the environment, PS-PRM needs to update the perception scores of its nodes and potentially replan the path if necessary. To enable fast, end-to-end replanning, we create a GPU-parallelized pipeline that performs collision checking, constraint projection, forward kinematics, and neural surrogate model inference in batches on the GPU. As the robot executes its planned trajectory, it continuously monitors whether the object has moved. If movement is detected (e.g., by the onboard camera), PS-PRM invokes the parallelized pipeline to re-project the existing nodes and update their perception scores. This involves batch collision checking against the new environment state, constraint projection for any nodes violating newly introduced constraints, batched forward kinematics to update camera transforms, and neural surrogate model inference to recalculate the perception scores.
@@
Nodes that cannot be projected into valid configurations or that collide in the new environment are pruned from the roadmap along with their associated edges. After pruning and re-scoring, the planner re-evaluates the current trajectory by computing a new average perception score; if that score falls below a specified threshold, it triggers a replanning step to find a more suitable path from the robot’s current position to the goal region. If the remaining nodes fail to produce a safe path to the goal, PS-PRM partially rebuilds the roadmap by sampling additional nodes while retaining any valid ones, thus avoiding a complete reconstruction. This leverages each step based on their computational needs, while building a PRM can be performed in seconds, re-evaluating nodes’ collisions and scores can be done at higher frequencies of several times a second. Through this GPU-accelerated parallelization, PS-PRM rapidly adapts to dynamic changes in the environment and maintains both collision-free and perception-optimized trajectories with reduced computation time. This results in a significant reduction in overall computation time and energy usage, enabling deployment in dynamic environments as shown in Experimental Results section Sec. V.
@@ V Experimental Results
In this section, we evaluate our proposed methods in simulation and real-robot experiments using the Hello Robot’s Stretch $2$ [89], Fetch robot, and a UR5 robot mounted on a differential-drive base. We use PyBullet [90] for collision checking, ompl’s Python bindings [91] for planning, and Isaac Sim [92] for simulation and visualization. All experiments were conducted on an Intel i7-12700K CPU and an NVIDIA GeForce RTX4090 GPU. We consider both object detection/tracking task and human face detection task. For the robot’s camera parameters in simulation experiments, we use a resolution of 640 $\times$ 480 pixels, a field of view of 42.5 degrees, and a clipping range of 0.3 meters for the near plane and 6 meters for the far plane.
@@
TABLE I: Planning and perception results for the Stretch robot in a simulated home environment. The results are averaged over 2,500 different motion planning problems. “Preprocess Time” is the time spent (e.g., building the PRM and evaluating the perception scores) before planning. “Plan. Time” is the average planning time. “Total Time” is the overall time including preprocessing and planning. “Path Len.” is the average length of the path. “Succ. Rate” is motion planning’s success rate. “Det. Rate” is the object detection rate. “Conf. Score” is the average confidence score of detections. “Track Rate” is the percentage of the frames that the object is tracked.
@@ V-A Evaluated Methods
We implement four versions of our approach ps-prm, as described in Sec. IV, varying along two key dimensions: the perception model (line-of-sight vs. camera frustum) and whether a neural surrogate is used for scoring.
@@
Specifically, in our experiments, we use the object detection confidence score as the perception score for object detection and tracking tasks, as described in Sec. V-B, Sec. V-E, and Sec. V-F. For human detection tasks, such as those discussed in Sec. V-C and Sec. V-D, we use the confidence score of the human face detector RetinaFace [93] from the deepface [94] as the perception score.
@@
This model is trained on a dataset composed of a diverse set of commonly encountered objects and humans from the COCO dataset [95], enabling generalization across a wide range of realistic detection scenarios. We generate a dataset of perception scores as the objects or human are viewed from different camera transform in Issac Sim. Specifically, we sample a set of 1000 camera transforms where the object or human are in the field of view uniformly at random. For each camera transform, we render the images and evaluate the perceptions scores using a perception model with the relevant perception tasks. Then, we train our model on the collected dataset for 200 epochs with an Adam [96] optimizer, learning rate of $1\times 10^{-4}$, and batch size of 512.
V-B Benchmark Results
@@
Table I presents results across 2,500 planning problems, evaluating metrics such as planning time, path length, success rate. We use YOLOv9 [97] and DeepSORT [98] for real-time perception performance assessment. It is measured by per-frame object detection rate, average confidence score of object detection, and tracking rate.
@@
It should be noted that constructing the constrained PRM with perception-based edge scores, using approximately 3,000 nodes across different monitored objects and scenes (in this experiment, 5 different objects and 5 locations each), takes on average 341.5 seconds for PS-PRM-line and 291.7 seconds for PS-PRM-frustum. Once the PRM is built, subsequent planning is very fast, as demonstrated in Table I.
@@
Fig. 6 illustrates an example of robot trajectories and corresponding camera view images for three different methods: post, Manifold-frustum, and PRM-perception. Both the robot paths and camera snapshots are sampled uniformly along the full trajectory. The figure demonstrates that the method which directly integrates perception scores, PS-PRM-line, achieves significantly better object detection and tracking performance compared to the baseline methods. Notably, it selects a trajectory that stays slightly farther from the monitored object, resulting in a better viewing angle at the cost of a slightly longer path.
@@
Figure 6: This figure shows one example of the robot trajectory and corresponding camera view images for three different methods (post, Manifold-frustum, and PRM-perception) for the simulation experiment in Sec. V-B. Both the robot traces and camera images are sampled evenly along the robot’s complete path. The robot’s initial position is shown in full opacity, gradually fading as it moves towards the goal. The camera view images are displayed in sequential order from top left to bottom right. If a snapshot does not display a score, it indicates that the object was not detected, and the score is zero.
V-C Human Detection in a Nursing Setting
As shown in Fig. 7, we create a nursing environment with two nurses with masks and one patient. The robot is required to plan a path from the start to the goal position while aiming to continuously detect some key features of the nurses. In this experiment, we consider the face detection task, and we use deepface and RetinaFace for real-time face detection performance assessment. For this multi-person task, we use the midpoint between the two individuals as the monitoring point and the average of their face detection scores to construct the PRM and assess performance. In this experiment, we use the Fetch robot and plan the motion of its base, torso, and camera to perform the assigned task.

### v2

@@ IV-A Neural Surrogate Perception Score Models
To address these challenges, we propose a neural surrogate model inspired by [72, 49] that approximates perception scores without requiring explicit rendering. Alg. 2 outlines the full pipeline. Given a robot configuration and the target object label, the surrogate model predicts the expected perception performance (e.g., detection confidence for a cup). We first apply FK to map the configuration from $\mathcal{Q}$ to the robot’s camera pose. The neural model is then trained to map these camera transforms to perception scores obtained from labeled, photorealistic simulations. Specifically, we generate a dataset by rendering views of objects or humans from varied camera poses in simulation and evaluating each image with a perception model. The surrogate is implemented as a multilayer perceptron (MLP) [76] that takes the relative pose between camera and object, along with an object encoding, and outputs the corresponding perception score. Trained on this dataset, the surrogate enables fast and scalable evaluation during planning. Importantly, because the input is based on camera-to-object pose rather than robot-specific joint values, the model generalizes across robots and environments with similar perception setups. This robot-agnostic formulation is central to our approach: by operating in SE(3) camera space rather than robot configuration space, the surrogate can be reused across different platforms without retraining, as demonstrated in our experiments.
To address object inter-occlusion in cluttered indoor environments, we build upon techniques similar to those in [102, 27] by performing ray-casting towards surrounding objects. This step determines whether the object of interest is occluded by checking for intersections between each cast ray and the Axis-Aligned Bounding Boxes (AABB) of nearby objects. For a new robot configuration, we obtain an occlusion score in addition to the perception score estimated by the neural surrogate model. This occlusion score is computed as the ratio of rays that ultimately do not intersect the bounding box of the target object, an indicator that something else is blocking the view. If the occlusion score exceeds a chosen threshold, we set the perception score of the neural surrogate model to zero for that configuration. Fig. 4 shows a detailed overview of this occlusion checking step.
@@ IV-B1 Sampling with Visibility Constraints
For high-d o f robots, computing motion that considers the aforementioned object-centered models can be achieved using sampling-based motion planning (SBMP) [64] with manifold constraints [39] where planning happens on a subspace of the free configuration space defined by a manifold. Specifically, in the construction of our proposed planner shown in Alg. 1, ps-prm first constructs a PRM with states and transitions that satisfy the manifold constraints, using either line-of-sight or camera frustum. The PRM is defined $G=(V,E)$, where $V$ is the set of configurations and $E$ the set of edges connecting them. The neighbor nodes in the PRM are constrained to lie on the same manifold and are selected based on their geodesic proximity within a certain distance. To sample from the manifold, ps-prm first selects a robot configuration from $\mathcal{Q}_{free}$ and then applies the projection operator to map it onto the surface of $M$. This involves iteratively retracting the point towards a minimum of the constraint function and solving a linear system of equations at each step. Configurations that cannot be projected into valid in-manifold configurations are discarded, and the algorithm resamples until a valid configuration is found or a timeout is reached. This process ensures that the object is placed either within the line of sight or inside the camera frustum.
@@ IV-B2 PRM construction and scoring
After the PRM is constructed, ps-prm connects the start and goal to the roadmap and applies an A* search [29] with an admissible heuristic to find a path from $q_{start}$ to $\mathcal{Q}_{goal}$ for the robot to follow.
Next, to guarantee the optimality of the A* search, we show that our A* heuristic remains admissible under the modified edge-cost formulation. Since consistency implies admissibility [18], it is sufficient to prove that our heuristic is consistent. We define our A* heuristic for a PRM $G=(V,E)$ as a “hop”-based function that lower-bounds the remaining motion and perception costs:
$h(q)\;=\;H_{\min}\cdot c^{\min},$ (3)
where $H_{\min}(q)$ denotes the hop distance, i.e., the minimum number of edges required to reach the goal from the configuration $q$ obtained via a shortest-path search, and $c^{\min}$ represents the minimum edge cost over the entire roadmap:
$c^{\min}=\min_{(u,v)\in E}c(q_{u},q_{v}).$ (4)
For any edge $(u,v)\in E$, a feasible path from $u$ to the goal is to first take $(u,v)$ and then follow a shortest path from $v$ to the goal. Thus, we have:
$H_{\min}(q_{u})\;\leq\;1+H_{\min}(q_{v}).$ (5)
Therefore, our heuristic function $h(\cdot)$ satisfies:
$h(q_{u})$ $=H_{\min}(q_{u})\cdot c^{\min}$
$\leq(1+H_{\min}(q_{v}))\cdot c^{\min}$
$\leq c(q_{u},q_{v})+h(q_{v})$ Eqs. 3 and 4
which is exactly the condition of a consistent heuristic. Thus, we show that our A* heuristic remains consistent and admissible under the modified edge cost formulation.
This formulation only requires nonnegativity of edge costs, without assuming a specific form such as Euclidean distance for motion cost. It applies broadly, e.g., when $c_{m}$ is defined as the trajectory length, energy, or control effort, and $c_{p}$ is a non-negative normalized neural perception score. In this paper, we demonstrate experiments on object detection and tracking, as well as human face detection and activity recognition, though the approach can be generalized to other perception tasks.
Since our problem is formulated as a sampling-based motion planning problem, probabilistic completeness [18] is an important property. We therefore show that probabilistic completeness still holds after introducing our modified edge-cost formulation. Specifically, we show that our perception-weighted PRM formulation does not invalidate the standard PRM completeness. This holds because the reweighting affects only the edge costs, not the set of feasible configurations or connections. Assume a PRM constructs a graph $G=(V,E)$ over samples from the collision-free configuration space $Q_{\mathrm{free}}$. The PRM completeness proof on Theorem 7.4.1 from [18], does not depend on edge cost function, only on samples and connections. Let $\sigma$ denote any feasible path with clearance $\delta>0$. With a probability exponentially increasing with the number of sampled states, PRM will include vertices $\{q_{1},\ldots,q_{n}\}$ such that consecutive vertices are within a local planner’s connection radius and $\sigma=(q_{1},\ldots,q_{n})$ will be found. Since our composite edge cost $c(u,v)$ shown in Eq. 2 is nonnegative and finite for all $(u,v)\in E$, the reweighting does not eliminate any edges or feasible connections. Consequently, PS-PRM remains probabilistically complete: if a feasible path exists in $Q_{\mathrm{free}}$, the probability that PS-PRM finds a finite-cost path approaches one as the number of samples increases.
@@ IV-C GPU-Parallelized Replanning Pipeline
The proposed perception scoring methods are highly parallelizable. Since the perception scores are computed using a neural surrogate model specifically optimized for GPU inference, we are motivated to implement the entire planning pipeline on the GPU. For collision checking, PS-PRM evaluates the validity of sampled configurations with hardware acceleration: instead of checking each obstacle individually, the algorithm performs batched collision checking on all obstacles simultaneously. Furthermore, both FK and surrogate model inference are performed in batches and executed in the GPU. This node-level parallelization accelerates roadmap construction and perception-scoring and enables real-time replanning in high-dimensional spaces. PS-PRM can therefore operate effectively in dynamic environments where the object of interest, goals or obstacles may move during execution. In such cases, the planner updates node perception scores in real time and replans the path as needed to adapt to the new environment. Fig. 5 shows a detailed overview of the GPU-accelerated process.
As the robot executes its planned trajectory, it continuously monitors whether the object has moved. If movement is detected (e.g., by the onboard camera), PS-PRM invokes the parallelized pipeline to re-project the existing nodes and update their perception scores.
@@
Nodes that cannot be projected into valid configurations or that collide in the new environment are pruned from the roadmap along with their associated edges. After pruning and re-scoring, the planner re-evaluates the current trajectory by computing a new average perception score; if that score falls below a specified threshold, it triggers a replanning step to find a more suitable path from the robot’s current position to the goal region. If the remaining nodes fail to produce a safe path to the goal, PS-PRM partially rebuilds the roadmap by sampling additional nodes while retaining any valid ones, thus avoiding a complete reconstruction. This leverages each step based on their computational needs, while building a PRM can be performed in seconds, re-evaluating nodes’ collisions and scores can be done at higher frequencies of several times a second. Through this GPU-accelerated parallelization, PS-PRM rapidly adapts to dynamic changes in the environment and maintains both collision-free and perception-optimized trajectories with reduced computation time.
In this work, we primarily consider small to medium target or obstacle motions, where changes in the environment are gradual relative to the planner’s replanning frequency. Under these conditions, the proposed planner can update trajectories online by replanning fast enough to adapt to dynamic targets, obstacles, and goals. Specifically, Sec. V-E demonstrates the feasibility of our system. Large or abrupt target motions are not the primary focus of this work. When the target dynamics exceed the planner’s replanning capability (e.g., rapid, discontinuous, or adversarial movements), the problem becomes fundamentally one of high-frequency tracking or reactive control rather than perception-constrained motion planning, which is outside the scope of this paper.
@@ V Experimental Results
In this section, we evaluate our proposed methods in simulation and real-robot experiments using the Hello Robot’s Stretch $2$ [37], Fetch robot, and a UR5. We use PyBullet [19] for collision checking, ompl’s Python bindings [87] for planning, and Isaac Sim [63] for simulation and visualization. All experiments were conducted on an Intel i7-12700K CPU and an NVIDIA GeForce RTX4090 GPU. We consider both object detection/tracking, human face detection and activity recognition as perception tasks. For the robot’s camera parameters in simulation experiments, we use a resolution of 640 $\times$ 480 pixels, a field of view of 42.5 degrees, and a clipping range of 0.3 meters for the near plane and 10 meters for the far plane.
@@
TABLE II: Planning and perception results for the Stretch robot in a simulated home environment. The results are averaged over 2,500 different motion planning problems. “Preprocess Time” is the time spent (e.g., building the PRM and evaluating the perception scores) before planning. “Plan. Time” is the average planning time. “Total Time” is the overall time including preprocessing and planning. “Path Len.” is the average length of the path. “Succ. Rate” is motion planning’s success rate. “Det. Rate” is the object detection rate. “Conf. Score” is the average confidence score of detections. “Track Rate” is the percentage of the frames that the object is tracked.
@@ V-A Evaluated Methods
All four baseline methods are tree-based planners derived from RRT-Connect, and therefore they do not require preprocessing time. We implement four versions of our approach ps-prm, as described in Sec. IV, varying along two key dimensions: the perception model (line-of-sight vs. camera frustum) and whether a neural surrogate is used for scoring.
@@
Specifically, in our experiments, we use the object detection confidence score as the perception score for object detection and tracking tasks, as described in Sec. V-C, Sec. V-F, and Sec. V-G. For human tracking tasks, such as those discussed in Sec. V-D and Sec. V-E, we use the confidence score of the human face detector RetinaFace [21] from Deepface [78] as the perception score.
@@
This model is trained on a dataset composed of a diverse set of commonly encountered objects and humans from the COCO dataset [47], enabling generalization across a wide range of realistic detection scenarios. We generate a dataset of perception scores as the objects or human are viewed from different camera transform in Issac Sim. Specifically, we sample a set of 50000 camera transforms where the object or human are in the field of view uniformly at random. For each camera transform, we render the images and evaluate the perceptions scores using a perception model with the relevant perception tasks. Then, we train our model on the collected dataset for 200 epochs with an Adam [38] optimizer, learning rate of $1\times 10^{-4}$, and batch size of 512.
For the dynamic experiments in Sec. V-E, all evaluations are conducted using the GPU implementation, as the CPU-based version is not able to run and replan at a rate for dynamic scenarios. This highlights the necessity of fast replanning when targets, obstacles, or goals are moving, and demonstrates the practical advantage of GPU parallelism in enabling real-time perception-constrained planning. By leveraging massively parallel FK and perception evaluation, the GPU implementation achieves the replanning frequency required to respond to environmental changes, making dynamic operation feasible in practice.
Estimation Error Ranking Detection Classification
Method RMSE $\downarrow$ MAE $\downarrow$ Spearman $\rho$ $\uparrow$ Acc. $\uparrow$ Prec. $\uparrow$ Rec. $\uparrow$ F1 $\uparrow$
Surrogate Model 0.154 0.085 0.641 93.1% 74.4% 65.5% 69.7%
TABLE III: Benchmark results of the neural surrogate model “Estimation Error“ (RMSE, MAE) measures the deviation of the predicted confidence score from ground truth. “Ranking“ (Spearman’s $\rho$) measures the model’s ability to correctly order camera poses by quality. “Detection Classification“ metrics (Accuracy, Precision, Recall, F1) evaluate the binary visibility prediction with a confidence threshold.
V-B Perception Scoring Benchmark
We evaluate the neural surrogate model from three complementary perspectives: confidence estimation accuracy, ranking consistency, and binary detection performance. First, estimation quality is assessed using Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) between the predicted confidence scores and YOLO-based ground truth, capturing absolute prediction errors. Second, since the planner primarily relies on relative perception quality rather than exact confidence values, we evaluate ranking consistency using Spearman’s rank correlation coefficient $\rho$, which measures how well the model preserves the ordering of camera poses by perceptual quality. Finally, we evaluate binary perception score prediction using standard classification metrics (accuracy, precision, recall, and F1 score) with a confidence threshold of $0.75$.
As shown in Table III, the surrogate model achieves low estimation error, indicating accurate confidence prediction, while maintaining a strong positive rank correlation, demonstrating its effectiveness in preserving perceptual ordering across camera viewpoints. Moreover, the high classification accuracy and balanced precision–recall performance confirm that the model reliably distinguishes higher scored configurations. Together, these results indicate that the surrogate model provides both accurate and decision-relevant perception estimates, making it well-suited for perception-guided motion planning.
The ray-tracing accuracy is also tested, comparing the visibility rate given by our proposed system using a single ray test, giving a score of 1 if the ray hits it’s target AABB and 0 otherwise. An environment with a perception target (a human) similar to Fig. 4 is evaluated with a wall, a thin wall (partial occlusion) and a fence-like wall (so target is visible through). The test was evaluated against a YOLO-based boolean detection with a $0.75$ confidence threshold. The single-ray test proved to be a conservative estimator in the fence and thin wall environments, reporting an $90.1\%$ and $85.1\%$ visibility rate compared to YOLO’s $97.0\%$ and $95.0\%$, as the ray is easily intercepted by thin geometric features despite the target being actually visible through gaps. Conversely, it slightly overestimated visibility in solid occlusion ($54.5\%$ estimated vs. $49.0\%$ ground truth detection), likely due specific angles and target-specific inaccuracies, such as a human being more easily distinguishable if the face or silhouette are visible. Overall, the single-ray test approximated the ground truth within a $10\%$ margin, which allowed our use for it as a computationally efficient occlusion estimator.
V-C Planning Benchmark Results
@@
Table II presents results across 2,500 planning problems, evaluating metrics such as planning time, path length, success rate. We use YOLOv9 [95] and DeepSORT [99] for real-time perception performance assessment. It is measured by per-frame object detection rate, average confidence score of object detection, and tracking rate.
@@
It should be noted that constructing the constrained PRM with perception-based edge scores, using approximately 3,000 nodes across different monitored objects and scenes (in this experiment, 5 different objects and 5 locations each), takes on average 341.5 seconds for PS-PRM-line and 291.7 seconds for PS-PRM-frustum. Once the PRM is built, subsequent planning is very fast, as demonstrated in Table II.
@@
Fig. 6 illustrates an example of robot trajectories and corresponding camera view images for three different methods: Post, Manifold-frustum, and PS-PRM-line. Both the robot paths and camera snapshots are sampled uniformly along the full trajectory. The figure demonstrates that the method which directly integrates perception scores, PS-PRM-line, achieves significantly better object detection and tracking performance compared to the baseline methods. Notably, it selects a trajectory that stays slightly farther from the monitored object, resulting in a better viewing angle at the cost of a slightly longer path. We notice that the confidence score is sometimes higher when detecting the back of the screen than the front. This occurs because the back of the monitor is dark and high-contrast relative to the background, whereas the front is lighter and contains glossy or reflective regions. These properties can make the back side easier for the detector to localize, leading to a higher confidence score. In general, detection confidence produced by a perception model can vary due to multiple factors, including the distribution of training data, color and texture contrast, lighting conditions, surface reflectance, and viewpoint-dependent features. As the learned perception score is trained across a wide range of environments, viewpoints, and object configurations, these artifacts are mitigated and smoothed.
@@
Figure 6: This figure shows one example of the robot trajectory and corresponding camera view images for three different methods (post, Manifold-frustum, and PRM-perception) for the simulation experiment in Sec. V-C. Both the robot traces and camera images are sampled evenly along the robot’s complete path. The robot’s initial position is shown in full opacity, gradually fading as it moves towards the goal. The camera view images are displayed in sequential order from top left to bottom right. If a snapshot does not display a score, it indicates that the object was not detected, and the score is zero.
V-D Human Tracking in a Nursing Setting
As shown in Fig. 7, we create a nursing environment with two nurses with masks and one patient. The robot is required to plan a path from the start to the goal position while aiming to continuously detect some key features of the nurses. In this experiment, we consider the face detection task, and we use RetinaFace [21] for real-time face detection performance assessment. For this multi-person task, we use the midpoint between the two individuals as the monitoring point and the average of their face detection scores to construct the PRM and assess performance. In this experiment, we use the Fetch robot and plan the motion of its base, torso, and camera to perform the assigned task.